# A Systematic Survey on Large Language Models for Evolutionary Optimization: From Modeling to Solving

### v2

@@ VI-B Natural Sciences
In the natural sciences, including chemistry [209], biology [210], and physics [211], many fundamental problems can be cast as complex optimization tasks. These problems are often characterized by high-dimensional search spaces, implicit domain knowledge, and computationally expensive simulations. The reasoning and generative capabilities of LLMs provide a novel approach to addressing these challenges.
In chemistry, LLM-assisted optimization has been widely explored in molecular discovery [209]. Most studies emphasize low-level methods that embed LLMs into EAs. For example, Wang et al. [212] integrated LLMs enhanced with chemical knowledge to perform crossover and mutation, guiding molecular design more effectively. Guevorguian et al. [213] replaced genetic operators with fine-tuned LLMs and introduced a prophetic model to avoid stagnation, later extending this idea to multi-objective optimization for molecular discovery [214]. Beyond operator-level assistance, some works explored surrogate modeling, such as Nguyen et al. [137], who applied LLMs to mitigate data scarcity in molecular optimization.
Similar to chemistry, biological applications often adopt low-level paradigms. The LLM-GA framework [210] uses LLMs for population initialization, generating high-quality mutant libraries for enzyme design. Tran et al. [215] and Wang et al. [216] employed LLMs as intelligent mutation or crossover operators for protein design. Beyond operator-level integration, Chen et al. [217, 120] developed a two-level optimizer that iteratively modifies biological sequences, while Lv et al. [218] proposed a workflow spanning modeling and solving, translating natural language descriptions into protein function predictions and optimizing sequences iteratively.
Applications in physics are more diverse, spanning fluid dynamics [211, 219], semiconductor physics [220], and statistical physics [221, 222]. At the modeling level, Du et al. [221] used GPT-3.5 to derive partial differential equations from data, while Zhang et al. [211] applied LLMs to turbulence closure modeling. Ma et al. [223] proposed a holistic workflow that integrates both modeling and solving. At the solving level, Li et al. [220] incorporated LLMs into optimization workflows to generate and debug code for optimizing laser parameters, while Zhang et al. [219] directly used LLMs as optimizers to generate parameter combinations for fluid dynamics.
@@ VI-C Engineering and Industry
In engineering and industry, LLMs are increasingly applied to complex optimization problems. Representative domains include wireless networks and communications [122], industrial design [170], and edge computing and scheduling [162], with wireless networks and industrial design being the most active areas.
In wireless networks, LLMs have been employed across the full workflow from modeling to solving. On the modeling side, Peng et al. proposed LLM-OptiRA [224], which automatically detects non-convex components in wireless resource allocation problems and reformulates them into solvable forms. Wen et al. [225] further introduced Retrieval-Augmented Generation (RAG) to incorporate external expert knowledge, improving modeling accuracy. On the solving side, Qiu et al. [122] directly used LLMs as optimizers for combinatorial tasks such as access point placement, with subsequent work extending this paradigm to resource allocation [226], power control [227], and multi-UAV deployment [228]. Li et al. [229] explored low-level assistance by embedding LLMs into a multi-objective optimization framework for ISAC systems in UAV networks. More recently, Hou et al. proposed the LHS framework [230], combining heuristic recommendation, enhanced initialization, and iterative refinement.
In industrial design, LLMs have been used for both algorithm configuration and generation. Ghose et al. [124] designed an LLM-based agent for chip design, enabling parameter tuning to optimize performance, power, and area. Jiang et al. [121] applied LLMs as optimizers for design structure matrices, capturing interdependencies within engineering systems. Yin et al. [170] explored algorithm generation in optical design, applying the LLaMEA framework [157] to discover novel algorithms for photonic structure design. These examples highlight how LLMs introduce fundamentally new approaches to long-standing industrial optimization problems.
Beyond networks and design, LLMs are also applied in edge computing and scheduling. For instance, Yatong et al. [162] leveraged LLMs for task scheduling in edge server environments, demonstrating their utility in distributed system optimization.
@@ VII-A Bridging Modeling and Solving
Although LLMs have been successfully applied to both optimization modeling and solving, these domains remain largely disjoint. Frameworks such as OptiMUS [71, 79] and ORLM [27] begin to connect the pipeline: OptiMUS employs an agent-based conductor for modeling and programming, while ORLM fine-tunes LLMs to generate mathematical models and solver code. Yet in all cases, the crucial solving step still depends on external, traditional solvers. This confines LLMs to the front-end, without incorporating advances in LLM-assisted solving such as adaptive configuration, selection, or algorithm generation.
@@ VII-B Transitioning from Static to Dynamic Methods
Current research in LLMs for optimization spans tasks such as optimization modeling, algorithm configuration, selection, and generation. Yet most approaches remain static, relying on fixed prompts, predefined operators, or single-shot algorithm designs. By contrast, the reinforcement learning community has developed rich paradigms for dynamic configuration [47, 231], selection [232], and generation [17], where algorithm behavior adapts continuously to real-time optimization features. For example, dynamic selection methods can exploit the complementary strengths of multiple algorithms through adaptive scheduling, while dynamic generation can synthesize novel operators tailored to evolving search states.
In LLM-based optimization, only a handful of works explore dynamic behavior, such as interactive modeling [80] or adaptive parameter control [139, 145]. Critically, dynamic algorithm selection and generation remain almost untouched. We envision a decisive shift toward methods that not only adapt in real time but also enable the self-evolution of algorithms: frameworks in which LLMs can continually refine, recombine, and reinvent optimization strategies based on feedback from the search process itself. Such self-evolving systems would transform LLMs from passive assistants into adaptive entities capable of lifelong learning, bridging optimization with open-ended algorithmic innovation.
@@ VII-C Toward an Agentic Ecosystem for Optimization
The role of LLMs in optimization should ultimately be conceived not as isolated engines, but as autonomous agents within a broader agentic ecosystem for optimization. In such systems, LLMs function as specialized agents collaborating with others (e.g., human experts, domain-specific solvers, and simulators) through communication, negotiation, and coordination. Early studies already suggest this potential: LLMs can capture tacit expertise, such as driver knowledge in routing [233], domain constraints in scheduling [234], and user preferences in traveling [235], making it actionable within optimization workflows [70, 236, 80, 202]. Similarly, agentic interfaces have enhanced accessibility and trust, where LLM-based agents diagnose infeasibility in natural language [237], explain the details of decision making [238, 239], or interpret optimization processes and outcomes for end users [240, 241, 242].
@@ VIII Conclusion
This survey has presented a comprehensive and systematic review of the emerging field of LLMs for optimization. As an interdisciplinary synthesis of LLMs and optimization methods, this field is reshaping the foundations of optimization research, bringing new perspectives to long-standing challenges and inspiring a series of novel paradigms. By structuring the discussion from modeling to solving, we have aimed to clarify how LLMs can participate in, and ultimately transform, the entire optimization workflow.

### v3

@@ VI-B Natural Sciences
In the natural sciences, including chemistry [wang2025survey], biology [nana2025integrating], and physics [zhang2025autoturb], many fundamental problems can be cast as complex optimization tasks. These problems are often characterized by high-dimensional search spaces, implicit domain knowledge, and computationally expensive simulations. The reasoning and generative capabilities of LLMs provide a novel approach to addressing these challenges.
In chemistry, LLM-assisted optimization has been widely explored in molecular discovery [wang2025survey]. Most studies emphasize low-level methods that embed LLMs into EAs. For example, Wang et al. [wang2024efficient] integrated LLMs enhanced with chemical knowledge to perform crossover and mutation, guiding molecular design more effectively. Guevorguian et al. [guevorguian2024small] replaced genetic operators with fine-tuned LLMs and introduced a prophetic model to avoid stagnation, later extending this idea to multi-objective optimization for molecular discovery [ran2025mollm]. Beyond operator-level assistance, some works explored surrogate modeling, such as Nguyen et al. [nguyen2024lico], who applied LLMs to mitigate data scarcity in molecular optimization.
Similar to chemistry, biological applications often adopt low-level paradigms. The LLM-GA framework [nana2025integrating] uses LLMs for population initialization, generating high-quality mutant libraries for enzyme design. Tran et al. [tran2024protein] and Wang et al. [wang2025large] employed LLMs as intelligent mutation or crossover operators for protein design. Beyond operator-level integration, Chen et al. [chen2024llms, chen2024generalists] developed a two-level optimizer that iteratively modifies biological sequences, while Lv et al. [lv2025prollama] proposed a workflow spanning modeling and solving, translating natural language descriptions into protein function predictions and optimizing sequences iteratively.
Applications in physics are more diverse, spanning fluid dynamics [zhang2025autoturb, zhang2025using], semiconductor physics [li2023english], and statistical physics [du2024large, sun2025integrating]. At the modeling level, Du et al. [du2024large] used GPT-3.5 to derive partial differential equations from data, while Zhang et al. [zhang2025autoturb] applied LLMs to turbulence closure modeling. Ma et al. [ma2024llm] proposed a holistic workflow that integrates both modeling and solving. At the solving level, Li et al. [li2023english] incorporated LLMs into optimization workflows to generate and debug code for optimizing laser parameters, while Zhang et al. [zhang2025using] directly used LLMs as optimizers to generate parameter combinations for fluid dynamics.
@@ VI-C Engineering and Industry
In engineering and industry, LLMs are increasingly applied to complex optimization problems. Representative domains include wireless networks and communications [qiu2024large], industrial design [yin2025optimizing], and edge computing and scheduling [yatong2024ts], with wireless networks and industrial design being the most active areas.
In wireless networks, LLMs have been employed across the full workflow from modeling to solving. On the modeling side, Peng et al. proposed LLM-OptiRA [peng2025llm], which automatically detects non-convex components in wireless resource allocation problems and reformulates them into solvable forms. Wen et al. [wen2025hybridrag] further introduced Retrieval-Augmented Generation (RAG) to incorporate external expert knowledge, improving modeling accuracy. On the solving side, Qiu et al. [qiu2024large] directly used LLMs as optimizers for combinatorial tasks such as access point placement, with subsequent work extending this paradigm to resource allocation [lee2024llm], power control [zhou2024large], and multi-UAV deployment [wang2025multi]. Li et al. [li2025large] explored low-level assistance by embedding LLMs into a multi-objective optimization framework for ISAC systems in UAV networks. More recently, Hou et al. proposed the LHS framework [hou2024wireless], combining heuristic recommendation, enhanced initialization, and iterative refinement.
In industrial design, LLMs have been used for both algorithm configuration and generation. Ghose et al. [ghose2025orfs] designed an LLM-based agent for chip design, enabling parameter tuning to optimize performance, power, and area. Jiang et al. [jiang2024large] applied LLMs as optimizers for design structure matrices, capturing interdependencies within engineering systems. Yin et al. [yin2025optimizing] explored algorithm generation in optical design, applying the LLaMEA framework [van2024llamea] to discover novel algorithms for photonic structure design. These examples highlight how LLMs introduce fundamentally new approaches to long-standing industrial optimization problems.
Beyond networks and design, LLMs are also applied in edge computing and scheduling. For instance, Yatong et al. [yatong2024ts] leveraged LLMs for task scheduling in edge server environments, demonstrating their utility in distributed system optimization.
@@ VII-A Bridging Modeling and Solving
Although LLMs have been successfully applied to both optimization modeling and solving, these domains remain largely disjoint. Frameworks such as OptiMUS [ahmaditeshnizi2024optimus, ahmaditeshnizi2024optimus03] and ORLM [huang2025orlm] begin to connect the pipeline: OptiMUS employs an agent-based conductor for modeling and programming, while ORLM fine-tunes LLMs to generate mathematical models and solver code. Yet in all cases, the crucial solving step still depends on external, traditional solvers. This confines LLMs to the front-end, without incorporating advances in LLM-assisted solving such as adaptive configuration, selection, or algorithm generation.
@@ VII-B Transitioning from Static to Dynamic Methods
Current research in LLMs for optimization spans tasks such as optimization modeling, algorithm configuration, selection, and generation. Yet most approaches remain static, relying on pre-specified configurations and offline designs rather than adapting algorithm behavior in response to real-time optimization signals or feedback. By contrast, the reinforcement learning community has developed rich paradigms for dynamic configuration [tian2022deep, sharma2019deep], selection [guo2024deep], and generation [chen2024symbol], where algorithm behavior adapts continuously to real-time optimization features. For example, dynamic selection methods can exploit the complementary strengths of multiple algorithms through adaptive scheduling, while dynamic generation can synthesize novel operators tailored to evolving search states.
In LLM-based optimization, only a handful of works explore dynamic behavior, such as interactive modeling [lawless2024want] or adaptive parameter control [custode2024investigation, zhang2025laos]. Critically, dynamic algorithm selection and generation remain almost untouched. We envision a decisive shift toward methods that not only adapt in real time but also enable the self-evolution of algorithms: frameworks in which LLMs can continually refine, recombine, and reinvent optimization strategies based on feedback from the search process itself. Such self-evolving systems would transform LLMs from passive assistants into adaptive entities capable of lifelong learning, bridging optimization with open-ended algorithmic innovation.
@@ VII-C Toward an Agentic Ecosystem for Optimization
The role of LLMs in optimization should ultimately be conceived not as isolated engines, but as autonomous agents within a broader agentic ecosystem for optimization. In such systems, LLMs function as specialized agents collaborating with others (e.g., human experts, domain-specific solvers, and simulators) through communication, negotiation, and coordination. Early studies already suggest this potential: LLMs can capture tacit expertise, such as driver knowledge in routing [liu2023can], domain constraints in scheduling [jobson2024investigating], and user preferences in traveling [li2024research], making it actionable within optimization workflows [xiao2023chain, li2024stride, lawless2024want, wang2025agenttts]. Similarly, agentic interfaces have enhanced accessibility and trust, where LLM-based agents diagnose infeasibility in natural language [chen2024diagnosing], explain the details of decision making [kikuta2024routeexplainer, wasserkrug2024large], or interpret optimization processes and outcomes for end users [singh2024enhancing, chacon2024large, maddigan2024explaining].
@@ VIII Conclusion
This survey has presented a structured review of the emerging field of LLMs for optimization. As an interdisciplinary synthesis of LLMs and optimization methods, this field is reshaping the foundations of optimization research, bringing new perspectives to long-standing challenges and inspiring a series of novel paradigms. By structuring the discussion from modeling to solving, we have aimed to clarify how LLMs can participate in, and ultimately transform, the entire optimization workflow.
@@
This section serves to consolidate and present the detailed taxonomies of the literature reviewed in the main body of the survey, ensuring thoroughness and ease of reference for the reader. Specifically, Table II provides a comprehensive breakdown of works focused on LLMs for optimization modeling, detailing their specific objectives and the methodological classification of their approach. Furthermore, Table III systematically catalogs the research concerning LLMs for optimization solving, classified according to the specific role of the LLM within the overall optimization workflow (e.g., as a low-level operator or a high-level generator).
Method Venue Type Technical Summary Prompt-based Methods Ner4OPT [dakle2023ner4opt] CPAIOR, 2023 Two-stage Fine-tune models for named entity recognition by incorporating traditional NLP methods. AOMG [almonacid2023towards] GECCO, 2023 Direct Directly leverage LLMs to generate mathematical models. HG 2.0 [tsouros2023holy] arXiv, 2023 Two-stage Embed LLMs within the two-stage framework. AMGPT [li2023synthesizing] arXiv, 2023 Two-stage Employ fine-tuned models for constraint classification. CoE [xiao2023chain] ICLR, 2023 Multi-agent Construct dynamic reasoning chains using 11 expert agents. OptiMUS [ahmaditeshnizi2024optimus] ICML, 2023 Multi-agent Develop a conductor agent to coordinate modeling, programming, and evaluation processes. MAMS [mostajabdaveh2024optimization] INFOR, 2024 Multi-agent Develop inter-agent cross-validation to replace solver-dependent verification. EC [jin2024democratizing] SGC, 2024 Two-stage Apply the two-stage framework to energy management systems. CAFA [deng2024cafa] NeurIPS, 2024 Two-stage Enhance modeling performance through code-based formalization. MAMO [huang2024llms] NAACL, 2024 Two-stage Develop the MAMO benchmark with ordinary differential equation extensions. NL2OR [li2025abstract] arXiv, 2024 Two-stage Predefine abstract structural constraints to regulate the LLM outputs. TRIP-PAL [de2024trip] arXiv, 2024 Two-stage Apply the two-stage framework to travel planning. OptLLM [zhang2024solving] arXiv, 2024 Interactive Support both single-input and interactive-input modes. OptiMUS-0.3 [ahmaditeshnizi2024optimus03] arXiv, 2024 Multi-agent Introduce self-corrective prompts and structure-aware modeling based on OptiMUS. MeetMate [lawless2024want] TiiS, 2024 Interactive Develop an interactive system for user input processing with five selectable task options. VL [gomez2025values] CUI, 2025 Interactive Convert user priorities into optimization constraints during dialog interactions. LLM-MCTS [astorga2024autoformulation] arXiv, 2025 Two-stage Conduct hierarchical Monte Carlo tree search over the hypothesis space. EquivaMap [zhai2025equivamap] arXiv, 2025 Two-stage Generate variable mapping functions via LLMs with lightweight verification. MAP [talebi2025large] arXiv, 2025 Multi-agent Deploy multiple independent reviewers to evaluate modeling results. ORMind [wang2025ormind] arXiv, 2025 Multi-agent Replace the conductor agent with structured, predictable workflows. LEAN-LLM-OPT [liang2025llm] ssrn, 2025 Multi-agent Integrate RAG to enhance problem classification, and construct instances with query operations. Learning-based Methods LM4OPT [ahmed2024lm4opt] INFOR, 2024 Fine-tuning Progressively fine-tune models on the NL4OPT dataset. ORLM [huang2025orlm] OR, 2024 Data synthesis Synthesize data through expansion and augmentation and fine-tune open-source models. ReSocratic [yang2024optibench] ICLR, 2024 Data synthesis Propose inverse data synthesis methodology and construct the OPTIBENCH benchmark. LLMOPT [jiang2024llmopt] ICLR, 2024 Fine-tuning Introduce model alignment and self-correction mechanisms to mitigate hallucination phenomena. BPP-Search [wang2024bpp] arXiv, 2024 Data synthesis Solve the problem of missing detailed in data synthesis. OptMATH [lu2025optmath] arXiv, 2025 Data synthesis Develop a scalable bidirectional data synthesis approach. LLMBO [amarasinghe2023ai] arXiv, 2025 Data synthesis Propose a method for fine-tuning cost-effective LLMs to tackle specific business challenges. SIRL [chen2025solver] arXiv, 2025 Fine-tuning Integrate external optimization solvers as verifiable reward validators for reinforcement learning. Step-Opt [wu2025step] arXiv, 2025 Data synthesis Increase problem complexity through an iterative problem generation approach. DPLM [zhou2025auto] arXiv, 2025 Data synthesis Combine the diversity of forward generation and the reliability of inverse generation. OptiTrust [lima2025toward] arXiv, 2025 Data synthesis Develop a verifiable synthetic data generation pipeline. StepORLM [zhou2025steporlm] arXiv, 2025 Fine-tuning Introduce generative process supervision and a co-evolutionary loop to refine reasoning trajectories.
Method Venue Type Technical Summary LLMs as Optimizers OPRO [yang2023large] ICLR, 2023 Prompt-based Iteratively refine through optimization trajectories and problem descriptions. toLLM [guo2023towards] KDD, 2023 Prompt-based Design four canonical tasks to evaluate the performance boundaries of LLMs. EvoLLM [lange2024large] GECCO, 2024 Prompt-based Replace traditional optimization trajectories with candidate solution quality ranking. MLLMO [huang2025multimodal] MCII, 2024 Prompt-based Utilize MLLMs to jointly process problem descriptions and map visualizations for CVRPs. POM [li2024pretrained] NeurIPS, 2024 Learning-based Pre-train a general-purpose, zero-shot black-box optimization foundation model. OPTO [cheng2024trace] NeurIPS, 2024 Prompt-based Replace traditional optimization trajectories with rich, structured execution traces. VRMA [elhenawy2024visual] arXiv, 2024 Prompt-based Utilize MLLMs to process 2D planar point distribution maps as input. ROPRO [zhang2024revisiting] ACL, 2024 Prompt-based Discuss the model dependency of OPRO and identify its limitations on small-scale models. BBOLLM [huang2024exploring] arXiv, 2024 Prompt-based Conduct evaluation of LLMs on both discrete and continuous black-box optimization problems. LLMS [abgaryan2024llms] arXiv, 2024 Learning-based Fine-tune LLMs with instruction-solution pairs for scheduling problems. ECLIPSE [jiang2024optimizable] NAACL, 2024 Prompt-based Apply iterative optimization to jailbreaking attack strategies. LLOME [chen2024generalists] arXiv, 2024 Learning-based Integrate preference learning to train LLMs for satisfying complex biophysical constraints. LLMDSM [jiang2024large] arXiv, 2024 Prompt-based Apply iterative optimization to design structure matrix sequencing. LMCO [qiu2024large] WCL, 2024 Prompt-based Apply iterative optimization to wireless network design. MGSCO [zhao2025bridging] arXiv, 2025 Prompt-based Process visual representations of abstract graph structures using MLLMs. ORFS [ghose2025orfs] arXiv, 2025 Prompt-based Apply iterative optimization to automated parameter tuning in chip design. Low-level LLMs for Optmization Algorithms LMX [meyerson2024language] TELO, 2023 Operators Leverage LLMs as intelligent operators for textual genome crossover and recombination. GPT-NAS [yu2023gpt] arXiv, 2023 Initialization Utilize LLMs for NAS initialization with prior knowledge. LMEA [liu2024large] CEC, 2023 Operators Employ LLMs as crossover, mutation, and selection operators to guide EAs. LLM-PP [jawahar2023llm] arXiv, 2023 Initialization Utilize LLMs as performance predictors to assist initialization processes. LMOEA [liu2025large] EMO, 2023 Operators Empower MOEA/D with LLMs through zero-shot prompting as search operators in multi-objective optimization. OFPLLM [de2023optimized] AI, 2023 Initialization Assist non-expert users in initializing financial plans. LEO [brahmachary2025large] Neucom, 2024 Operators Utilize LLMs to guide individuals from dual pools for exploration and exploitation. LLM-GA [teukam2024integrating] BiB, 2024 Initialization Initialize genetic algorithms with LLMs to generate high-quality mutant pools for enzyme design. GE [morris2024llm] GECCO, 2024 Operators Enhance the creativity and diversity of the LLM by introducing different role-based prompts for NAS. LLMAES [wang2024largeicic] ICIC, 2024 Operators Generate minimal solutions within populations using LLMs to reduce interaction costs. LLMTES [kramer2024large] arXiv, 2024 Configuration Apply LLM-based feedback loops to sequentially optimize evolution strategies through static tuning. LAEA [hao2024large] SWEVO, 2024 Evaluation Transform model-assisted selection tasks into classification and regression problems. LICO [nguyen2024lico] arXiv, 2024 Evaluation Apply LLMs as surrogate models in molecular science applications. LLMSM [rios2024large] CAI, 2024 Evaluation Develop a collaborative framework using LLMs for model selection and training in engineering optimization. LLMCES [custode2024investigation] GECCO, 2024 Configuration Implement step-size control for evolution strategies using OPRO-like mechanisms. AKFLLM [huang2024advancing] arXiv, 2024 Operators Utilize LLMs to assist mutation and other generative stages in evolutionary multi-task optimization. LLMAMO [liu2024largemulti] arXiv, 2024 Operators Invoke LLMs for elite solution generation only upon insufficient population improvement. LTC [kramer2024llama] ESANN, 2024 Configuration Apply LLM-based feedback loops to sequentially optimize CMA-ES through dynamic control. LLMEVO [zhao2025can] arXiv, 2025 Operators Validate LLM effectiveness in selection, crossover, and mutation phases while noting limitations in initialization. LLM-GE [yu2025llm] GECCO, 2025 Operators Utilize LLMs as crossover and mutation operators to optimize YOLO architectures for object detection. LAOS [zhang2025laos] GECCO, 2025 Configuration Utilize state features to replace optimization trajectories for guiding LLM-based operator selection. PAIR [ali2025pair] arXiv, 2025 Operators Focus on utilizing LLMs as selection operators to enhance LMEA. LLMMS [zhang2025large] arXiv, 2025 Evaluation Utilize LLMs as meta-Surrogates to effectively facilitate cross-task knowledge sharing through token sequence representations. LMPSO [shinohara2025large] arXiv, 2025 Operators Simulate PSO evolutionary processes using LLMs and refine LMEA for specific algorithms. LLM-SAEA [xie2025large] arXiv, 2025 Evaluation Dynamically select appropriate proxy models and infill sampling criteria using LLMs. High-level LLMs for Optmization Algorithms LLMGM [pluhacek2023leveraging] GECCO, 2023 Single-step Decompose and recombine six existing metaheuristic algorithms to generate novel hybrid algorithms in single rounds. FunSearch [romera2024mathematical] Nature, 2023 Iterative Generate novel code snippets via LLMs and guide EAs to search through function space. AEL [liu2023algorithm] arXiv, 2023 Iterative Enhance FunSearch by integrating heuristic principles, demonstrating significant improvements on TSP instances. AS-LLM [wu2023large] IJCAI, 2023 Selection Automatically extract high-dimensional algorithm features from code and descriptive text to facilitate algorithm selection. EoH [liu2024evolution] ICML, 2024 Iterative Implement heuristic-code coevolution via preset prompt pairs for algorithm crossover and mutation. ReEvo [ye2024reevo] NeurIPS, 2024 Iterative Utilize LLMs as hyper-heuristic algorithms with reflective evolution mechanisms to guide search processes. AutoSAT [sun2024autosat] arXiv, 2024 Iterative Employ multiple heuristic strategies to guide LLMs in algorithm generation. LLaMEA [van2024llamea] TEVC, 2024 Iterative Utilize LLMs to guide EAs in mutation, selection, and other operations for generating advanced heuristics. L-AutoDA [guo2024autoda] GECCO, 2024 Iterative Apply AEL to generate adversarial attack algorithms in cybersecurity. MOELLM [huang2025autonomous] TEVC, 2024 Iterative Introduce robust testing modules and dynamic selection strategies to generate multi-objective optimization algorithms. LLMEPS [zhang2024understanding] PPSN, 2024 Iterative Develop a baseline algorithm for automated algorithm design using EoH and ReEvo methodologies. EvolCAF [yao2024evolve] PPSN, 2024 Iterative Apply EoH to evolve and generate cost-aware acquisition functions for Bayesian optimization. TS-EoH [yatong2024ts] arXiv, 2024 Iterative Apply EoH to generate algorithms for edge server task scheduling. MEoH [yao2025multi] AAAI, 2024 Iterative Extend EoH to multi-objective versions, incorporating other metrics for comprehensive algorithm evaluation. LLaMEA-HPO [van2024loop] TELO, 2024 Iterative Integrate hyperparameter optimization into LLaMEA’s iterative cycle to enhance overall efficiency. CMLLM [yin2025controlling] EvoApps, 2024 Iterative Develop dynamic prompt mechanisms to control LLM mutation for enhancing LLaMEA performance. HSEvo [dat2025hsevo] AAAI, 2024 Iterative Combine harmony search with GAs and optimize algorithm portfolio quality/diversity through diversity metrics. LLM4AD [liu2024llm4ad] arXiv, 2024 Iterative Develop an EoH-based platform to accelerate heuristic algorithm design. MCTS-AHD [zheng2025monte] arXiv, 2025 Iterative Employ MCTS to explore heuristic spaces, organizing generated heuristics into tree structures. CEG [van2025code] GECCO, 2025 Iterative Develop methodologies to analyze LLM-generated code and its evolutionary patterns during optimization processes. OPSLAD [yin2025optimizing] GECCO, 2025 Iterative Apply LLaMEA to photonic structure optimization in industrial design. BLADE [van2025blade] GECCO, 2025 Iterative Establish a standardized and reproducible evaluation framework for LLM-driven algorithm generation methods. AutoHD [ling2025complex] arXiv, 2025 Iterative Propose an LLM-guided heuristic discovery strategy for complex planning tasks. LAMA [li2025llm] TEVC, 2025 Iterative Propose an automatic memetic algorithm enhanced by a heuristic designed with the assistance of LLMs. LAS [liu2025fitness] arXiv, 2025 Iterative Analyze fitness landscapes of LLM-assisted algorithm search processes. LLaMEA-BO [li2025llamea] arXiv, 2025 Iterative Apply LLaMEA to automatically generate Bayesian optimization algorithms. InstSpecHH [zhang2025llm] arXiv, 2025 Selection Leverage LLMs’ semantic comprehension and reasoning capabilities for context-aware algorithm selection. BSALMD [van2025behaviour] arXiv, 2025 Iterative Conduct behavioral space analysis to examine algorithmic evolution trajectories and behaviors. FLAAD [liu2025fine] arXiv, 2025 Iterative Fine-tune LLMs with diversity-aware ranked sampling for algorithm generation. EoH-S [liu2025eoh] arXiv, 2025 Iterative Extend EoH to evolve and generate complementary algorithm portfolios.